How many millimoles of KSCN is in 750 mL of 3.25×10-3 M solution?


1
Expert's answer
2022-02-15T16:47:04-0500

Solution:

Moles of solute = Molarity × Liters of solution

Hence,

Moles of KSCN = (3.25×10-3 M) × (750 mL) × (1 L / 1000 mL) = 0.0024375 mol KSCN

Millimoles of KSCN = (0.0024375 mol) × (1000 mmol / 1 mol) = 2.4375 mmol = 2.44 mmol


Answer: 2.44 millimoles of KSCN